28. Regulatory transcription factors (the proteins themselves)

A. are trans-acting factors.

B. are cis-acting elements.

C. are promoters.

D. are the same in all cells regardless of tissue type.

E. are encoded by centromeres.

29. The process that leads to development haploid gamete is best described as ________.

A. segregation

B. independent assortment

C. mendelian inheritance

D. dominance

30. A cross between two individuals with different phenotypes that resulted in approximately 50% of

each type of offspring would indicate the cross was ________.

A.

true breeding dominant to recessive

B.

a heterozygous dominant crossed to a heterozygous recessive

C.

a heterozygous dominant crossed to a homozygous recessive

D.

a homozygous recessive crossed to a heterozygous recessive

33. Mendel's Law of Segregation is supported by a ________ testcross ratio.

A. 1:1

B. 3:1

C. 1:2:1

D. 1:2

34. Assuming complete dominance, a phenotypic ratio of ________ is expected from a monohybrid sib

or self-cross

A. 1:2

B. 3:1

C. 3:2

D. 1:2

35. Germ cell mutations ________

A. are of no consequence to the next generation

B. affect somatic cell function in the current generation

C. are usually found in every cell of the next generation

D. are only seen in females

36. Mutations that arise in nature, from no particular artificial agent, are called ________.

A. induced mutations

B. spontaneous mutations

C. chromosomal losses
Answer #1

28) transcription factors are proteins that bind to the promoters of the gene and activate the transcription of the gene, some transcription factors are specific to the tissue type which activates the expression of a particular gene. transcription factors are trans-acting factors that activate the expression of the genes, so the answer is A. are trans-acting factors.

29) the process by which haploid gametes are produced from diploid cells is called meiosis. alleles segregate each other during gametogenesis, so the answer is A) segregation.

30) let the alleles be D-dominant, d-recessive

Dd * dd

D d
d Dd dd

percentage of Dd progenies= (number of Dd/total number)100=1/2100=50%.

percentage of dd progenies= (number of dd/total number)100=1/2100=50%.

so each phenotype of the offspring are obtained in equal proportion so the answer is

C.a heterozygous dominant crossed to a homozygous recessive

33) let the alleles be D-dominant, d-recessive

then the test cross is

Dd * dd

D d
d Dd dd

the ratio is Dd:dd=1:1

so the answer is A) 1:1